OpenStudy (anonymous):

For the investment with interest compounded annually, find the final balance using the formula A = P(1 + r)t $5000 for 5 years at 3.2% A. $5160.00 B. $5852.86 C. $20,037.32 D. $25,800.00

OpenStudy (anonymous):

plz help!

OpenStudy (anonymous):

i got d.

OpenStudy (anonymous):

hurry

OpenStudy (anonymous):

5800

OpenStudy (anonymous):

@BeautifulSoulღ

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

i got b or d

OpenStudy (anonymous):

@crazysingh @cheergurl_99 @coolsday @cutegirl @Cutefriendzoned

OpenStudy (anonymous):

@CaseyCarns

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

@annas @arabpride

OpenStudy (anonymous):

i will give medal and fan

OpenStudy (anonymous):

answer is B.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

r/100 = 3.2/100 = 0.032 A = P(1+r)^t A = 5000(1+0.032)^5 A = 5000(1.032)^5 solve for A.

OpenStudy (anonymous):

umm..

OpenStudy (anonymous):

not a

OpenStudy (anonymous):

\[5000\times(1+\frac{ 3.2 }{ 100 })^{5} = 5000\times (1.032)^{5}=5852.86\]

OpenStudy (anonymous):

i put b

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

What is the final balance for an account starting with $2000 at 3.5% interest compounded annually for 3 years? A = P(1 + r)t A. $2217.44 B. $4290.75 C. $6210.00 D. $8100.00

OpenStudy (anonymous):

this one also

OpenStudy (anonymous):

A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

\[2000\times(1+\frac{ 3.5 }{ 100 })^{3}= 2217.43\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

City Bank offers 4% simple interest on deposits. County Bank offers 3% interest compounded annually. Derek deposits $5000 in City Bank and Lupe deposits $5000 in County Bank. What is the difference in final balances for Derek and Lupe at the end of 6 years? A. $229.74 B. $300.00 C. $356.33 D. $426.60

OpenStudy (anonymous):

@VrindaG

OpenStudy (anonymous):

help

OpenStudy (anonymous):

vrinda plz help

OpenStudy (anonymous):

in city bank total becomes \[5000+ (5000\times.04\times6)=6200\] in country bank total becomes \[5000\times(1+\frac{ 3 }{ 100 })^{6}=5970.261\] so the difference is 229.74 so the answer is A

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

Lindsay invested $4500 at 4% interest compounded annually. How much interest will she earn in 10 years? A. $180.00 B. $187.27 C. $1800.00 D. $2161.10

OpenStudy (anonymous):

last 1

OpenStudy (anonymous):

plz help!

OpenStudy (anonymous):

i got C.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

wait let me check ..

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

answer is D.

OpenStudy (anonymous):

i got c r u sure ur right!?

OpenStudy (anonymous):

\[4500\times(1+\frac{ 4 }{ 100 })^{10}-4500= 6661.099-4500 = 2161.099\]

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

is that your school homework ??

OpenStudy (anonymous):

ya it is d thx!

OpenStudy (anonymous):

thx for all the helpp bye!

OpenStudy (anonymous):

bye good night :)

OpenStudy (anonymous):

i got a 80% by the way!yay

OpenStudy (anonymous):

good night what its only 1:39 PM here in Nashville,TN

OpenStudy (anonymous):

are those online problems ??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

but its 12:10 am here

OpenStudy (anonymous):

which site are you studying ??

OpenStudy (anonymous):

im in online school its sunday but i feel like doing work

OpenStudy (anonymous):

K12 im in TNVA

OpenStudy (anonymous):

is it online? wow

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i just wanna be ahead in math!

OpenStudy (anonymous):

im 11 in 6th

OpenStudy (anonymous):

you can ask me problems and learn. I will help you.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

http://www.k12.com/tnva#.UyXwa6iSy-0

OpenStudy (anonymous):

Is the number prime or composite? 29 A. prime B. composite

OpenStudy (anonymous):

i think a

OpenStudy (anonymous):

it is prime .. since it can be divided by 1 and 29 only.

OpenStudy (anonymous):

State whether the number is prime or composite. 57 A. prime B. composite

OpenStudy (anonymous):

prime i bet

OpenStudy (anonymous):

19 x 3 =57 .. so composite

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you doing the another test.. ??

OpenStudy (anonymous):

Which example supports the given conjecture? The sum of three odd numbers is an odd number. A. 3 + 5 + 7 = 15 B. 1 + 2 + 3 = 6 C. 2 + 4 + 8 = 14 D. 4 + 5 + 6 = 15

OpenStudy (anonymous):

a

OpenStudy (anonymous):

a i knew it

OpenStudy (anonymous):

are you doing the another test.. ??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Which number supports the given conjecture? If a number is divisible by both 2 and 3, then the number is also divisible by 6. A. 16 B. 23 C. 27 D. 36 D right

OpenStudy (anonymous):

@crazysingh

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Which number supports the given conjecture? If a number is divisible by both 2 and 3, then the number is also divisible by 6. A. 16 B. 23 C. 27 D. 36

OpenStudy (anonymous):

its D. right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wrong one

OpenStudy (anonymous):

What are the first five terms of the pattern with the formula 21n? The first five terms are , , , , and .

OpenStudy (anonymous):

21, 42, 63, 84, 105

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

put n = 1, 2, 3, 4, 5

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

and multiply

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

sorry idk i cant help you im terrible at math